# Overdense fireworks in GOODS-N: Unveiling a record number of massive dusty star forming galaxies at z5.2 with the N2CLS

### v1

@@ 5.2 Source properties
The multi-wavelength SEDs for these 10 sources are shown in B25. Using CIGALE (see Sect. 3.2), we derive their properties that are summarised in Table 4. Our sample is characterised by high V-band attenuation (AV), SFR and stellar mass, with median values of AV=2.93, SFR=681 M⊙ yr-1 and log=11. These values are comparable to those reported for the three ultra-massive galaxies in Xiao et al. (2024)
@@
However, they are systematically much higher than those found in the larger DSFG sample from JWST (36 galaxies) analyzed in the same study.
@@
Combining the derived SFR and , and the known redshift, we apply the Tacconi et al. (2020) scaling relations to derive the molecular gas depletion time scale, , and mass, of these N2GN sources. Figure 2 locates the sources in the vs. SFR plane and compares their position to the main sequence (MS) loci defined by Speagle et al. (2014) and Popesso et al. (2023). The Tacconi et al. (2020) scalings are based on the distance of the sources from the Speagle et al. (2014) MS, which does not include the bending at large , appearing clearly instead in Popesso et al. (2023).
@@
The N2GN z5.2 galaxies lie significantly above the star-forming main sequence (MS) as defined by Popesso et al. (2023), indicating a strongly bursty mode of star formation. Although they appear closer to the MS when adopting the Speagle et al. (2014) parametrization, eight galaxies still lie above the 1 envelope of the MS.
@@
The median obtained for these galaxies is of 0.19 Gyr, with individual values spanning over the range between and Gyr. These short time scales (or equivalently very high star formation efficiency, SFE, yr-1) confirm the intense ongoing starbursts activity that powers their FIR and mm emission. Despite their short depletion timescales, the starbursts are still operating over periods significantly longer than their dynamical timescales, suggesting that the activity may be sustained through multiple bursts or extended starburst phases. The fact that these galaxies reside in a confirmed overdense environment likely provides both the fuel and favorable conditions to maintain such prolonged star formation.
@@
Given their short gas depletion timescales, these DSFGs may transition into a quiescent phase by , assuming a constant SFE, complete molecular gas consumption, and no external gas replenishment. This evolutionary pathway would be consistent with the emergence of massive quiescent galaxies observed in other overdense environments at (e.g., Carnall et al. 2024; Kakimoto et al. 2024; Tanaka et al. 2024).
Figure 2 also displays the locations of the DSFGs identified by Xiao et al. (2024), selected based on their JWST colors mag, which suggest significant dust attenuation. Notably, these JWST-selected galaxies lie systematically below the star-forming main sequence. This offset may stem from underestimated SFRs and/or overestimated M⋆. In the absence of far-infrared or (sub-)millimeter constraints (”No FIR” case), UV-based SFRs, corrected for dust attenuation, frequently fail to capture the bulk of star formation in dusty systems. GOODS-ALMA studies (Elbaz et al. 2018; Xiao et al. 2023) report typical ratios of total to dust-corrected UV SFRs on the order of 8. Conversely, Pacifici et al. (2023) find that including FIR data in SED fitting yield systematically lower SFRs compared to fits that omit such data. To investigate this further, we performed SED fitting on a subset of our N2CLS DSFGs using only optical-to-NIR photometry (m). Using BAGPIPES (Carnall et al. 2018), we modeled the UV-to-NIR SEDs of six sources with available HST+JWST photometry (N2GN_1_01, 06, 13, 23, 33, and 43). For N2GN_1_01 (S2) and N2GN_1_23 (S3), we confirm the results of Xiao et al. (2024), with SFRs significantly underestimated. However, this is not the case for N2GN_1_33 and N2GN_1_43, where higher SFRs are recovered. To further probe this discrepancy, we repeated the analysis using the CIGALE code. The inferred SFRs from the No IR fits differ notably from those obtained with the full UV-to-(sub)mm coverage, with differences reaching up to a factor of 30. SFRs are overestimated in the No IR scenario, except for two sources. Importantly, the stellar masses derived in the No IR case are systematically overestimated, with offsets ranging from a factor of 1.1 up to 7.5. These comparisons underscore that, even with the unprecedented depth of JWST, (sub-)mm data remain essential to accurately determine SFRs, M⋆ (and AV) in dusty galaxies. However, such data require very deep observations. Indeed, none of the 25 DSFGs in the GOODS-N field reported by Xiao et al. (2024) (except for the brightest sources N2GN_1_01, 06, and 23) are detected at in N2CLS, which already reaches the confusion limit in GOODS-N (Ponthieu et al. 2025).
@@
We finally examined the rest-frame optical morphologies of our sources using JWST/NIRCam imaging (Fig. 7). Most galaxies appear extended or clumpy, with no strong evidence for point-like or compact morphologies—only one galaxy in the sample appears marginally compact. Several galaxies show irregular or asymmetric light distributions, suggestive of ongoing or recent dynamical activity, such as minor mergers, interactions, or internal disk instabilities. These features are broadly consistent with recent JWST results for high-redshift, dust-obscured galaxies (e.g., McKinney et al. 2025), which reveal a wide range of morphologies, from clumpy and irregular to disk-like systems, among massive DSFGs.
@@ 6 Stellar baryon budget and the efficiency of star formation
Recent observations have unveiled a substantial population of massive galaxies in the early universe, challenging existing models of galaxy formation. Studies leveraging data from the JWST have identified galaxies at redshifts to with stellar masses approaching 1011 M⊙. For instance, utilizing JWST/MIRI data, Wang et al. (2024) found high densities of massive galaxies at , implying elevated stellar baryon efficiencies . Specifically, they report that the most massive galaxies at 8 require , compared to , for typical low-redshift galaxies.
@@
Similarly, Xiao et al. (2024) examined 36 massive, dust-obscured galaxies with spectroscopic redshifts between and from the JWST FRESCO survey. They found that three ultra-massive galaxies (log M⋆/M11.0, among them S2N2GN_1_01 and S3N2GN_1_23) require an exceptional of approximately 50%, which is two to three times higher than the efficiencies observed in galaxies at later epochs.
@@
Further, Gentile et al. (2024) reported on two massive and dusty starburst galaxies at z discovered in the COSMOS-Web survey. Their analysis indicates of approximately 0.25 and 0.5, respectively.
@@
In numerical simulations, is found to increase significantly with both stellar mass and redshift, indicating that, at fixed mass, halos at (very) high redshifts convert gas into stars more efficiently than their lower-redshift counterparts (e.g., Ceverino et al. 2024).
@@
All these results imply that certain populations of high-redshift galaxies may experience very efficient star formation at early times.
@@
Building upon these studies, we have computed the stellar baryon conversion efficiency for our sample of galaxies at , following the methodology adopted in Boylan-Kolchin (2023); Xiao et al. (2024).
@@
Using the cumulative halo mass function, we compute the comoving number density of halos above a given halo mass threshold, at redshift . The maximum stellar mass () that is statistically allowed within an observed volume is given by , where represents the (integrated) efficiency with which baryons are converted into stars and =0.156 is the cosmic baryon fraction from Planck Collaboration et al. (2020). We show on Fig. 3 our 10 galaxies in the stellar mass vs. redshift plane. The lines correspond to different number densities:
@@
3.710-6 Mpc-3 (i.e. one source in the cosmic volume explored by N2CLS, assuming z=5.20.5 ; with and displayed as blue-dotted and grey-dashed lines, respectively) and 3.910-12 Mpc-3 (i.e. a single galaxy in the whole sky; with displayed as a continuous grey line).
Our analysis reveals notably high values (0.2 for all but N2GN_1_06 and 61), indicating that a significant fraction of baryonic mass in these galaxies has been converted into stars.
Two galaxies (NN2GN_1_20 and 33), identified through photometric redshifts, lie above the model expectations for the GOODS-N survey volume. This discrepancy could arise from an overestimation of their photometric redshifts, assumptions related to the IMF (for example, a top-heavy IMF could yield lower stellar masses by dex, resulting in slightly reduced values of ) or from the presence of an unusually dense concentration of massive halos in this region of the sky.
@@
The elevated values of observed in our sample suggest that these galaxies are undergoing unusually efficient star formation at . Given their probable high metallicities and host halo masses exceeding , it is unlikely that cold gas inflows are the primary driver of this efficiency. Instead, the observed efficiency may reflect a temporary suppression of stellar feedback, as proposed in feedback-free starburst scenarios for higher-redshift galaxies with lower metallicities (e.g., Dekel et al. 2023; Li et al. 2024b), or merger-driven inflows that channel gas into central regions and ignite intense starbursts, or compaction events that funnel enriched gas into central regions, triggering intense, short-lived starbursts.
@@ 7.1 Contribution to the star formation rate density at
Recent studies have shown that a significant fraction of the cosmic star formation rate density (SFRD) at high redshift can be driven by rare, massive galaxies with high star formation efficiencies. For instance, Xiao et al. (2024) found that at , ultra-massive galaxies with can account for up to % of the total SFRD (from Madau & Dickinson 2014). These findings highlight the critical role of massive, DSFGs in early cosmic star formation, motivating a closer look at their contribution in overdense environments such as GOODS-N.
The combined cosmic SFRD from the five N2CLS galaxies with spectroscopic redshifts777Assuming an area of 160 arcmin2 and to compute the volume. is M⊙ yr-1 Mpc-3, representing a factor of 1.390.06 excess relative to the total SFRD reported by Madau & Dickinson (2014) at similar redshift. Including the additional contribution from galaxies with photometric redshifts increases the total SFRD to M⊙ yr-1 Mpc-3, corresponding to a factor of 4.270.16 higher than the cosmic average.
The values are shown in a figure in Appendix 6.
The elevated SFRD reflects the fact that our sample is embedded within a confirmed overdense environment, where star formation is significantly enhanced relative to typical field regions at this epoch.
More broadly, current research suggests that massive galaxy overdensities at high redshift may be relatively common (e.g., Kashino et al. 2023; Wang et al. 2023; Helton et al. 2024a; Brinch et al. 2024). In the GOODS-N field, Lin et al. (2025) estimate the fraction of the cosmic SFRD hosted within the protocluster environment at z5.2, based on the number density of H emitters. They find that approximately 55% of the SFRD at this redshift is associated with the protocluster, somewhat higher than the 30% predicted by simulations (Chiang et al. 2017) but in good agreement with recent results based on the FLAMINGO simulation suite (Lim et al. 2024). Within the same volume considered in Lin et al. (2025) (i.e., 62 arcmin2 and ), our four DSFGs yield a star formation rate density of [M⊙ yr-1 Mpc-3]) = . This value is 1.6 times higher than the estimate based on the 150 H emitters alone, suggesting that a substantial fraction of star formation activity is missed even in deep JWST surveys of highly overdense environments.
@@ 7.2 Environmental localization of N2CLS DSFGs in the GOODS-N overdensity
Recent JWST observations from the JADES and FRESCO programs have revealed that the overdensity surrounding HDF850.1 represents one of the richest environments identified to date at , with approximately 100 galaxies at =5.17–5.20 distributed across 10 substructures within a (15 cMpc)3 volume (Herard-Demanche et al. 2025; Sun et al. 2024). This region is particularly extreme, as 28% of all –6.0 H emitters detected across the GOODS-N field are concentrated within a narrow redshift slice of . The spatial extent of this protocluster is consistent with the typical size of such structures predicted by Chiang et al. (2017) using a set of N-body simulations and semi-analytic models. In addition, Calvi et al. (2023) identified the galaxy GN10 (N2GN_1_01) as residing in an overdense region based on photometric redshifts, a result that has now been confirmed spectroscopically by Sun et al. (2024).
We show the locations of the N2CLS galaxies within the overdensity in Fig. 4. Cosmological simulations suggest that at , approximately 90% of baryonic and dark matter resides in unvirialized, filamentary large-scale structures (e.g., Haider et al. 2016). Sun et al. (2024) interpret the 3D spatial distribution of group member galaxies as being consistent with three distinct filamentary structures. For simplicity, we compare the positions of our galaxies relative to the filaments identified in Sun et al. (2024), rather than to the 18 substructures detailed in Herard-Demanche et al. (2025). The –5.20 overdensity, traced by filament-1 and filament-2, spans the full extent of the GOODS-North FRESCO field (62 arcmin2). Within this region lie four N2CLS DSFGs with spectroscopic redshifts: N2GN_1_06 and N2GN_1_23 are located in filament-1, while the newly identified N2GN_1_13 and N2GN_1_61 fall within filament-2. In filament-3, N2GN_1_26 (photometric redshift) is found along with GN10, and N2GN_1_43 lies nearby. N2GN_1_57 may trace a distinct, possibly new, structure. We note that some previously confirmed spectroscopic members of the protocluster (Walter et al. 2012; Calvi et al. 2021) fall outside the JADES and FRESCO joint footprint, implying that the overdensity may exhibit more complex substructures than currently captured. The sources N2GN_1_20 and N2GN_1_33 are located farther from the core overdensity. Interestingly, Jiang et al. (2018) report a giant protocluster at occupying a volume of cMpc3, embedded within a larger overdense region spanning cMpc2 and extending 34 cMpc along the line of sight. Further investigation is needed to determine whether the overdensity extends toward the regions traced by N2GN_1_20 and N2GN_1_33.
The spatial distribution of N2CLS galaxies at z5.2, when viewed in two dimensions (RA, DEC), does not exhibit a recognizable filamentary pattern, unlike the tentative identification of a filament with NIKA2 in the GJ526 field (Lestrade et al. 2022).

### v2

@@ 5.2 Source properties
The multi-wavelength SEDs for these eight sources are shown in B25. Here, we used CIGALE to derive their properties. As in B25, we adopted the bc03 stellar models, with a chabrier2003 IMF and solar metallicity, the charlot2000 attenuation law, and the DL07 dust emission models, re-normalised as in draine2014. A delayed and truncated star formation history was used, with e-folding time of the main stellar population of 0.1, 0.5, 1, 5, 10 Gyr, and truncation ages of 50 or 100 Myr.
To assess the impact of photometric-redshift uncertainties on the derived physical parameters, we repeated the fits at z1 for the three sources with photometric redshifts. The additional uncertainties were estimated from the standard deviation of the resulting physical parameters and added in quadrature to the fitting errors returned by CIGALE at zphot.
@@
We summarise the results in Table 4. Our sample is characterised by high V-band attenuation (AV), SFR and stellar mass, with median values of AV=3.26, SFR=587 M⊙ yr-1 and log()=10.93. These values are comparable to those reported for the three ultra-massive galaxies in Xiao2024
@@
However, they are systematically much higher than those found in the larger DSFG sample from JWST (36 galaxies) analysed in the same study.
@@
Combining the derived SFR and M⋆, and the known redshift, we applied the tacconi2020 scaling relations to derive the molecular gas depletion timescale, , and mass, of these N2GN sources. Figure 2 locates the sources in the versus SFR plane and compares their position to the MS loci defined by speagle2014 and popesso2023. The tacconi2020 scalings are based on the distance of the sources from the speagle2014 MS, which does not include the bending at large , appearing clearly instead in popesso2023.
@@
The N2GN z5.2 galaxies lie significantly above the SFMS as defined by popesso2023, indicating a bursty mode of star formation. Although they appear closer to the SFMS when adopting the speagle2014 parametrisation, the eight galaxies still lie above the envelope of the relation. This result will need to be confirmed for such high-mass and high-SFR galaxies at such redshifts once the SFMS is better constrained with JWST data.
@@
The median obtained for these galaxies is of 0.18 Gyr, with individual values spanning over the range between and Gyr. These short timescales (or equivalently a very high star formation efficiency, yr-1) confirm the intense ongoing starbursts activity that powers their far-IR and millimetre emission. Despite their short depletion timescales, the starbursts are still operating over periods significantly longer than their dynamical timescales, suggesting that the activity is sustained through multiple bursts or extended starburst phases. The fact that these galaxies reside in a confirmed overdense environment likely provides both the fuel and favourable conditions to maintain such prolonged star formation.
@@
Given their short gas depletion timescales, these DSFGs may transition into a quiescent phase by , assuming a constant star formation efficiency, complete molecular gas consumption, and no external gas replenishment. This evolutionary pathway would be consistent with the emergence of massive quiescent galaxies observed in other overdense environments at (e.g. Carnall2024; Kakimoto2024; Tanaka2024).
@@
We finally examined the rest-frame optical morphologies of our sources using JWST/NIRCam imaging (Fig. 7). Most galaxies appear extended or clumpy, with no strong evidence of point-like or compact morphologies ; only one galaxy in the sample appears marginally compact. Several galaxies show irregular or asymmetric light distributions, suggestive of ongoing or recent dynamical activity, such as minor mergers, interactions, or internal disk instabilities. These features are broadly consistent with recent JWST results for high-redshift, dust-obscured galaxies (e.g. McKinney2025), which reveal a wide range of morphologies, from clumpy and irregular to disk-like systems, among massive DSFGs.
5.3 Impact of missing far-IR data on SFR and stellar mass estimates for our sample
We also display in Fig. 2 the locations of the DSFGs identified by Xiao2024, selected based on their JWST colours mag, which suggest significant dust attenuation. Notably, these JWST-selected galaxies lie systematically below the SFMS. This offset may stem from underestimated SFRs and/or overestimated M⋆. In the absence of far-IR or (sub-)millimetre constraints (‘No IR’ case), UV-based SFRs, corrected for dust attenuation, frequently fail to capture the bulk of star formation in dusty systems. GOODS-ALMA studies (Elbaz2018; xiao2023) report typical ratios of total to dust-corrected UV SFRs of the order of 8. Conversely, Pacifici2023 find that including far-IR data in SED fitting yield systematically lower SFRs compared to fits that omit such data. To investigate this further, we performed SED fitting on a subset of our N2CLS DSFGs using only optical to near-IR photometry (m). Using BAGPIPES (Carnall2018), we modelled the UV to near-IR SEDs of five sources with available HST+JWST photometry (N2GN_1_01, 06, 13, 23 and 43). For N2GN_1_01 (S2) and N2GN_1_23 (S3), we confirm the results of Xiao2024, with SFRs significantly underestimated. However, this is not the case for N2GN_1_43 where higher SFRs are recovered. To further probe this discrepancy, we repeated the analysis using the CIGALE code. The inferred SFRs from the No IR fits differ notably from those obtained with the full UV to (sub-)millimetre coverage, with differences reaching up to a factor of 20. SFRs are overestimated in the No IR scenario, except for two sources. Importantly, the stellar masses derived in the No IR case are systematically overestimated, with offsets ranging from a factor of 1.2 up to 7.5. These comparisons underscore that, even with the unprecedented depth of JWST, (sub-)millimetre data remain essential to accurately determining SFRs and M⋆ (and AV) in dusty galaxies. However, such data require very deep observations. Indeed, none of the 25 DSFGs in the GOODS-N field reported by Xiao2024 are detected at in N2CLS, which already nearly reaches the confusion limit in GOODS-N (Ponthieu2025).
@@ 6 Stellar baryon budget and the efficiency of star formation
Recent observations have revealed a substantial population of massive galaxies in the early Universe, implying star-formation efficiencies that exceed the highest values inferred from lower-redshift studies. This may indicate a distinct and more intense mode of star formation during the earliest phases of galaxy evolution, which remains challenging for current models to reproduce (shen2025, and references therein).
@@
Studies leveraging data from the JWST have identified galaxies at redshifts to with stellar masses approaching 1011 M⊙. For instance, utilising JWST/Mid-infrared instrument (MIRI) data, Wang2025 found high densities of massive galaxies at , implying elevated stellar baryon efficiencies . Specifically, they report that the most massive galaxies at require , compared to , for typical low-redshift galaxies.
@@
Similarly, Xiao2024 examined 36 massive, dust-obscured galaxies with spectroscopic redshifts between and from the JWST FRESCO survey. They found that three ultra-massive galaxies (log(M⋆/M11.0, among them S2N2GN_1_01 and S3N2GN_1_23) require an exceptional of approximately 50%, which is two to three times higher than the efficiencies observed in galaxies at later epochs.
@@
Further, gentile2024 reported two massive, dusty starburst galaxies discovered in the COSMOS-Web survey, with star-formation efficiencies of 0.25 at =5.051 and 0.5 at =6.70.3, respectively.
@@
In numerical simulations, is found to increase significantly with both stellar mass and redshift, indicating that, at fixed mass, halos at (very) high redshifts convert gas into stars more efficiently than their lower-redshift counterparts (e.g. Ceverino2024).
@@
All these results imply that certain populations of high-redshift galaxies experience very efficient star formation at early times.
@@
Building upon these studies, we computed the stellar baryon conversion efficiency () for our sample of galaxies at , following the methodology adopted in Boylan-Kolchin2023 and Xiao2024.
@@
Using the cumulative halo mass function from tinker2008, we computed the comoving number density of halos above a given halo mass threshold, at redshift . The maximum stellar mass () that is statistically allowed within an observed volume is given by , where represents the (integrated) efficiency with which baryons are converted into stars and =0.156 is the cosmic baryon fraction from planck2020_cosmo. We show in Fig. 3 our 8 galaxies in the stellar mass versus redshift plane. The lines correspond to different number densities:
@@
2.210-6 Mpc-3 (i.e. one source in the cosmic volume explored by N2CLS, assuming z=5.20.5 ; with and displayed as blue-dotted and grey-dashed lines, respectively) and 2.410-12 Mpc-3 (i.e. a single galaxy in the whole sky; with displayed as a dotted grey line).
@@
Our analysis reveals notably high values (0.2 for all but N2GN_1_06 and 61), indicating that a significant fraction of baryonic mass in these galaxies has been converted into stars. The elevated values of observed in our sample suggest that these galaxies are undergoing unusually efficient star formation at . Given their probable high metallicities and host halo masses exceeding M⊙, it is unlikely that cold gas inflows are the primary driver of this efficiency. Instead, the observed efficiency may reflect merger-driven inflows that channel gas into central regions and ignite intense starbursts, or compaction events that funnel enriched gas into central regions, triggering intense, short-lived starbursts.
@@ 7.1 Contribution to the star formation rate density at
Recent studies have shown that a significant fraction of the cosmic SFRD at high redshifts can be driven by rare, massive galaxies with high star formation efficiencies. For instance, Xiao2024 found that at , ultra-massive galaxies with can account for up to % of the total SFRD (from Madau2014). These findings highlight the critical role of massive, DSFGs in early cosmic star formation, motivating a closer look at their contribution in overdense environments such as GOODS-N.
Based on the five N2CLS galaxies with spectroscopic redshifts, we derive a dust-obscured cosmic SFRD131313Assuming an area of 160 arcmin2 and to compute the volume. of M⊙ yr-1 Mpc-3. Including the additional contribution from sources with photometric redshifts raises the SFRD to M⊙ yr-1 Mpc-3. The quoted uncertainties combine Poisson errors and uncertainties on individual SFR estimates in quadrature.
Figure 4 illustrates the contribution of the N2CLS galaxies in the overdensity to the cosmic SFRD. For reference, we compared our measurements to the widely used compilation of Madau2014. The SFRD derived from spectroscopically confirmed galaxies, , is times higher than the total SFRD reported by MD14 at similar redshift. When including galaxies with photometric redshifts, the excess increases to a factor of . However, it is now well established that the MD14 compilation underestimates the SFRD at z4-6. ALMA studies (e.g. fujimoto2024; Traina2024; sun2025; liu2025) and JWST H luminosity function measurements (Covelo-Paz2025; Fu2025) both point to a systematically higher cosmic SFRD at these epochs. Using the updated estimate from fujimoto2024, the contrast between our overdensity and the global average decreases to about unity for and to 2 for . Restricting the comparison to the obscured component, our is already consistent with the levels reported by Traina2024 and sun2025, and rises substantially above them once photometric redshift members are included. Relative to the dust-obscured SFRD from Zavala2021 at similar redshifts, alone is more than five times higher. If all galaxies with photometric redshifts are spectroscopically confirmed, the elevated SFRD in our sample will reflect the fact that these sources reside in a confirmed overdense environment, where star formation activity is strongly boosted compared to average field regions at the same epoch.
Beyond our specific measurement, current research suggests that massive galaxy overdensities at high redshifts may be relatively common (e.g. Kashino2023; Wang2023; helton2024; Brinch2024) and may thus contribute significantly to the SFRD. In the GOODS-N field, Lin2025 estimate the fraction of the cosmic SFRD hosted within the protocluster environment at z5.2, based on the number density of H emitters. They find that approximately 55% of the SFRD at this redshift is associated with the protocluster, somewhat higher than the 30% predicted by simulations (Chiang2017) but in good agreement with recent results based on the FLAMINGO simulation suite (Lim2024). Within the same volume considered in Lin2025, our four DSFGs yield a SFR density of [M⊙ yr-1 Mpc-3]) = . This value is 1.6 times higher than the estimate based on the 150 H emitters alone, suggesting that a substantial fraction of star formation activity is missed even in deep JWST surveys of highly overdense environments.
@@ 7.2 Environmental localization of N2CLS DSFGs in the GOODS-N overdensity
Recent JWST observations from the JADES and FRESCO programmes have revealed that the overdensity surrounding HDF850.1 represents one of the richest environments identified to date at , with approximately 100 galaxies at =5.17–5.20 distributed across 10 substructures within a (15 cMpc)3 volume (Herard-Demanche2024; Sun2024). This region is particularly extreme, as 28% of all –6.0 H emitters detected across the GOODS-N field are concentrated within a narrow redshift slice of . The spatial extent of this protocluster is consistent with the typical size of such structures predicted by Chiang2017 using a set of N-body simulations and semi-analytic models. In addition, Calvi2023 identified the galaxy GN10 (N2GN_1_01) as residing in an overdense region based on photometric redshifts, a result that has now been confirmed spectroscopically by Sun2024.
We show the locations of the N2CLS galaxies within the overdensity in Fig. 5. Cosmological simulations suggest that at , approximately 90% of baryonic and dark matter resides in unvirialised, filamentary large-scale structures (e.g. Haider2016). Sun2024 interpret the 3D spatial distribution of group member galaxies as being consistent with three distinct filamentary structures. For simplicity, we compared the positions of our galaxies relative to the filaments identified in Sun2024, rather than to the 18 substructures detailed in Herard-Demanche2024. The –5.20 overdensity, traced by filament-1 and filament-2, spans the full extent of the GOODS-North FRESCO field (62 arcmin2). Within this region lie four N2CLS DSFGs with spectroscopic redshifts: N2GN_1_06 and N2GN_1_23 are located in filament-1, while the newly identified N2GN_1_13 and N2GN_1_61 fall within filament-2. In filament-3, N2GN_1_26 (photometric redshift) is found along with GN10, and N2GN_1_43 lies nearby. N2GN_1_57 may trace a distinct, possibly new, structure. We note that some previously confirmed spectroscopic members of the protocluster (walter2012; calvi2021) fall outside the JADES and FRESCO joint footprint, implying that the overdensity exhibits more complex substructures than currently captured.
We also show the location of N2GN_1_20, another N2CLS galaxy at (see Appendix D), which lies farther from the core overdensity. Interestingly, Jiang2018 report a giant protocluster at occupying a volume of cMpc3, embedded within a larger overdense region spanning cMpc2 and extending 34 cMpc along the line of sight. Further investigation is needed to determine whether the overdensity could extend towards the region traced by N2GN_1_20. Finally, we note that the spatial distribution of N2CLS galaxies at z5.2, when viewed in two dimensions (RA, Dec), does not exhibit a recognisable filamentary pattern, unlike the tentative identification of a filament with NIKA2 in the GJ526 field (Lestrade2022).